# Progressive Compositionality in Text-to-Image Generative Models

### v1

@@ 6 Conclusion
In this work, we present EvoGen, a curriculum contrastive framework to overcome the limitations of diffusion models in compositional text-to-image generation, such as incorrect attribute binding and object relationships. By leveraging a curated dataset of positive-negative image pairs and a multi-stage fine-tuning process, EvoGen progressively improves model performance, particularly in complex scenarios. Our experiments demonstrate the effectiveness of this method, paving the way for more robust and accurate generative models.
@@ A.2 Negative Text Prompts Generation
We apply in-context learning, everytim we prompt GPT-4 to generate negative cases, we give 5-10 example test prompts each time, and make sure the generation is not repetitive, within certain lengths.
@@
  • In Stage-II, we either swap the objects or attributes and let GPT-4 to validate the swapped text prompts. For complex cases, we generate negative text by asking GPT-4 to change the attributes/relationship/scenes.
  • In Stage-III, we carefully curate complicated examples with 3-6 objects, each object have 1-3 attributes, with negative prompts change attributes, actions and spatial relationships, scenes. And we prompt GPT-4 with such examples.
@@ A.3 VQA Assistance
Next, we illustrate how we prompt VQA to revise the caption when alignment scores of all candidate images are low. Given a generated image and a original text prompt, we prompt LLaVA “Given the original text prompt describing the image, identify any parts that inaccurately reflect the image. Then, generate a revised text prompt with correct descriptions, making minimal semantic changes.”. At the same time, we will give example of revised caption pairs to VQA for in-context learning.
@@ A.4 Data Statistics
Figure 9: Contrastive Images in ConPair
@@ Appendix B Training Implementation Details
We implement our apprpoach upon stable Diffuion v2.1 and we employ the pre-trained text encoder from the CLIP ViT-L/14 model. The VAE encoder is frozen during training. The resolution is 768, the batch size is 16, and the learning rate is 3e-5 with linear decay.
Appendix C Attn & Exct Benchmark Prompt Examples
@@
The benchmark protocol we follow comprises structured prompts ‘a [animalA] and a [animalB]’, ‘a [animal] and a [color][object]’, ‘a [colorA][objectA] and a [colorB][objectB]’ . Table 8 demonstrate the results of average CLIP similarities between text prompts and captions generated by BLIP for Stable Diffusion-based methods on this benchmark. EvoGen outperform those models on three categories.

### v2

@@ 6 Conclusion
In this work, we present EvoGen, a curriculum contrastive framework to overcome the limitations of diffusion models in compositional text-to-image generation, such as incorrect attribute binding and object relationships. By leveraging a curated dataset of positive-negative image pairs and a multi-stage fine-tuning process, EvoGen progressively improves model compositionality, particularly in complex scenarios. Our experiments demonstrate the effectiveness of this method, paving the way for more robust and accurate generative models.
@@ A.1 Text prompts generation
Note that when constructing our prompts, we consciously avoided using the same ones as those in T2I-Compbench, especially considering some prompts from T2I-CompBench are empirically difficult to generate aligned image (e.g., “a pentagonal warning sign and a pyramidal bookend” as shown in Figure 9), which are not well-suited for our dataset. We have filtered out similar prompts from our dataset using LLMs to identify uncommon combinations of objects and attributes.
Figure 9: Example image that is hard to generate to align the prompt from T2I-CompBench.
@@ A.2 Negative Text Prompts Generation
We apply in-context learning and prompt GPT-4 to generate negative cases, we give 5-10 example test prompts each time, and make sure the generation is not repetitive, within certain lengths.
@@
  • In Stage-II, we either swap the objects or attributes and let GPT-4 validate the swapped text prompts. For complex cases, we generate negative text by asking GPT-4 to change the attributes/relationships/scenes.
  • In Stage-III, we carefully curate complicated examples with 3-6 objects, each object has 1-3 attributes, with negative prompts change attributes, actions and spatial relationships, and scenes. We also prompt GPT-4 with such examples.
@@ A.3 VQA Assistance
Modifying Caption to Align Image.
Next, we illustrate how we prompt VQA to revise the caption when alignment scores of all candidate images are low. Given a generated image and an original text prompt, we prompt the VQA model with the following instruction:
Instruction: “Given the original text prompt describing the image, identify any parts that inaccurately reflect the image. Then, generate a revised text prompt with correct descriptions, making minimal semantic changes. Focusing on counting, color, shape, texture, scene, spatial relationship, and non-spatial relationship.”. At the same time, we will provide examples of revised captions for in-context learning.
For example, given the following image (Figure 10) and the original text prompt, the modified prompt generated by the VQA model is as follows:
Figure 10: Image applies reverse-alignment.
Original text prompt: Three puppies are playing on the sandy field on a sunny day, with two black ones walking toward a brown one.
Modified prompt: Four puppies are standing on a sandy field on a sunny day, with three black puppies and one brown puppy facing forward.
Note that the instruction ”Focusing on the counting, color, shape, texture, scene, spatial relationship, non-spatial relationship” plays a crucial role in guiding the VQA model to provide answers that accurately correspond to the specific attributes and categories we are interested in. Without this directive, the model may occasionally fail to generate precise captions that correctly describe the image.
@@ A.4 Data Statistics
Figure 11: Example contrastive Image pairs in ConPair
A.5 Comparison with Real Contrastive Dataset
To evaluate how our model would fare with a real hard-negative dataset, we include the results of fine-tuning our model with COLA (Ray et al., 2023), BISON (Hu et al., 2019) evaluated by T2I-CompBench in Table 7 (randomly sampled consistent number of samples across datasets).
Although COLA and BISON try to construct semantically hard-negative queries, the majority of the retrieved image pairs are quite different in practice, often introducing a lot of noisy objects/background elements in the real images, due to the nature of retrieval from an existing dataset. We hypothesize this makes it hard for the model to focus on specific attributes/relationships in compositionality. In addition, they don’t have complex prompts with multiple attributes and don’t involve action, or scene.
In contrast, our dataset ensures the generated image pairs are contrastive with minimal visual changes, enforcing the model to learn subtle differences in the pair, focusing on a certain category. To the best of our knowledge, no real contrastive image dataset only differs on minimal visual characteristics.
Dataset Color Shape Texture Spatial Non-Spatial Complex
COLA 62.20 48.98 53.73 15.21 30.87 33.15
BISON 59.49 49.36 48.77 14.64 31.25 32.91
Ours 71.04 54.57 72.34 21.76 33.08 42.52
Table 7: Performance of fine-tuning EvoGen on T2I-CompBench across different dataset.
Figure 12: Comparison with Real Contrastive Dataset: COLA and BISON.
A.6 Quality Control
Coverage of LLM-generated QA Pairs
We conducted human evaluations on Amazon Mechanical Turk (AMT). We sampled 1500 prompt-image pairs (about 10% of the dataset, proportionally across 3 stages) to perform the following user-study experiments. Each sample is annotated by 5 human annotators.
To analyze if the generated question-answer pairs by GPT-4 cover all the elements in the prompt, we performed a user study wherein for each question-prompt pair, the human subject is asked to answer if the question-set covers all the objects in the prompt. The interface is presented in Figure 13.
Figure 13: Interface for User Study: Coverage of LLM-generated QA Pairs
Empirically, we find about 96% of the questions generated by GPT-4 cover all the objects, 94% cover all the attributes/relationships.
Accuracy of Question-Answering of VQA Models
To analyze the accuracy of the VQA model’s answering results, we performed an additional user-study wherein for each question-image pair, the human subject is asked to answer the same question. The accuracy of the VQA model is then predicted using the human labels as ground truths. Results are displayed in Table 8.
Image Stage VQA Accuracy % Annotation Time / Image (s)
Stage-I 93.1% 8.7s
Stage-II 91.4% 15.3s
Stage-III 88.9% 22.6s
Table 8: VQA accuracy and annotation time for sampled images across different stages.
We observe that the VQA model is effective at measuring image-text alignment for the majority of questions even as the complexity of the text prompt increases, attesting the effectiveness of pipeline.
Alignment of Revised Caption with Images
To further validate the effectiveness of revising captions by VQA, we randomly sampled 500 images that are obtained by revising caption and performed an additional user-study for those samples that obtain low alignment score from VQA answering, but use the reverse-alignment strategy. Specifically, for each revised caption-image pair, the human subject is asked to answer how accurately the caption describes the image. The interface is presented in Figure 14. Note we have 5 annotators, each is assigned 100 caption-image pairs.
Figure 14: Interface for User Study: Alignment of Revised Caption with Images
Empirically, we found that 4% of the samples show that the revised caption similarly describes the image as the original caption. 94.6% of the samples show the revised caption better describes the image. Overall,with the following settings, the average rating of the alignment between revised caption and image is 4.66, attesting that revised caption accurately describes the image.
Similarity of Contrastive Image Pairs
We have 3 annotators in total, each annotator is assigned 2550 images (about 50% samples) to check if the positive and negative image pairs align with its text prompt and are similar with small visual changes on specific attributes/relationships. We filtered 647 images from the randomly selected 7650 images, which is 8.45%, attesting the quality of most images in the dataset.
@@ Appendix B Training Implementation Details
We implement our approach upon Stable Diffuion v2.1 and Stable Diffusion v3-medium. We employ the pre-trained text encoder from the CLIP ViT-L/14 model. The VAE encoder is frozen during training. The resolution is 768, the batch size is 16, and the learning rate is 3e-5 with linear decay.
Appendix C Quantitative Results
C.1 T2I-CompBench Evaluation Metrics
Following T2I-CompBench, we use DisentangledBLIP-VQA for color, shape, texture, UniDet for spatial, CLIP for non-spatial and 3-in-1 for complex categories.
C.2 Gen-AI Benchmark
We further evaluate EvoGen on the Gen-AI (Li et al., 2024a) benchmark. For a fair comparison with DALL-E 3, we finetune our model on Stable Diffusion v3 medium. As indicated in Table 9, EvoGen performs best on all the Advanced prompts, although it exhibits relatively weaker performance in some of the basic categories compared to DALL-E 3.
Basic Advanced
Method Attribute Scene Relation Avg Count Differ Compare Logical Avg
Spatial Action Part Negate Universal
SD v2.1 0.75 0.77 0.72 0.72 0.69 0.74 0.66 0.63 0.61 0.50 0.57 0.58
SD-XL Turbo 0.79 0.82 0.77 0.78 0.76 0.79 0.69 0.65 0.64 0.51 0.57 0.60
SD-XL 0.82 0.84 0.80 0.81 0.81 0.82 0.71 0.67 0.64 0.49 0.57 0.60
SD3-Medium 0.86 0.86 0.87 0.86 0.88 0.86 0.74 0.77 0.72 0.50 0.73 0.68
DALL-E 3 0.91 0.91 0.89 0.89 0.91 0.90 0.78 0.76 0.70 0.46 0.65 0.65
Table 9: Gen-AI Benchmark Results.
C.3 Attn & Exct Benchmark Prompt Examples
@@
The benchmark protocol we follow comprises structured prompts ‘a [animalA] and a [animalB]’, ‘a [animal] and a [color][object]’, ‘a [colorA][objectA] and a [colorB][objectB]’ . Table 10 demonstrate the results of average CLIP similarities between text prompts and captions generated by BLIP for Stable Diffusion-based methods on this benchmark. EvoGen outperforms those models in three categories.
@@ Appendix D Qualitative Results
Appendix E Related Work
With the rapid development of multimodal learning (Li et al., 2023; Liang et al., 2024b; a; Han et al., 2025) and image generation (Yu et al., 2024; Liu et al., 2024b; Weber et al., 2024; Peng et al., 2024b; Kim et al., 2025), understanding and addressing compositional challenges in text-to-image generative models has been a growing focus in the field (Thrush et al., 2022; Huang et al., 2023; Chefer et al., 2023; Peng et al., 2024c). In particular, Zarei et al. (2024) identifies key compositional challenges in text-to-image diffusion models and proposes strategies to enhance attribute binding and object relationships. Leveraging the power of large-language models (LLMs) for compositional generation is another area of active research (Drozdov et al., 2022; Mitra et al., 2024; Pasewark et al., 2024). For instance, Feng et al. (2023b) leverages large language models (LLMs) to generate visually coherent layouts and improve compositional reasoning in visual generation tasks.